# MDCure: A Scalable Pipeline for Multi-Document Instruction-Following

### v1

@@ Abstract
Multi-document (MD) processing is crucial for LLMs to handle real-world tasks such as summarization and question-answering across large sets of documents. While LLMs have improved at processing long inputs, MD contexts still present challenges, such as managing inter-document dependencies, redundancy, and incoherent structures. We introduce MDCure, a scalable and effective fine-tuning pipeline to enhance the MD capabilities of LLMs without the computational cost of pre-training or reliance on human annotated data. MDCure is based on generation of high-quality synthetic MD instruction data from sets of related articles via targeted prompts. We further introduce MDCureRM, a multi-objective reward model which filters generated data based on their training utility for MD settings. With MDCure, we fine-tune a variety of LLMs, from the FlanT5, Qwen2, and LLAMA3.1 model families, up to 70B parameters in size. Extensive evaluations on a wide range of MD and long-context benchmarks spanning various tasks show MDCure consistently improves performance over pre-trained baselines and over corresponding base models by up to 75.5%. Our code, datasets, and models are available at https://github.com/yale-nlp/MDCure.
@@
Gabrielle Kaili-May Liu1 Bowen Shi1 Avi Caciularu2 Idan Szpektor2  Arman Cohan1,3 1Yale University  2Google Research  3Allen Institute for AI kaili.liu@yale.edu     arman.cohan@yale.edu
@@ 1 Introduction
With the rapid expansion of bodies of online texts in domains of science, finance, education, news, and more, multi-document (MD) processing becomes a critical aspect of LLMs’ ability to understand, leverage, and reason over large quantities of text information, with applications such as (query-focused) multi-document summarization (MDS) (Fabbri et al., 2019; Xiao et al., 2022; Giorgi et al., 2023), multi-hop question-answering (MHQA) (Yang et al., 2018; Tu et al., 2019), and cross-document coreference resolution (Cattan et al., 2021b; Eirew et al., 2022).
While LLMs continue to upscale context size and improve performance on complex tasks, their ability to perform multi-document tasks is still limited (Hosseini et al., 2024; Lior et al., 2024; Wang et al., 2024d). MD processing presents unique challenges such as managing redundancy, high inter-document relatedness, contradictions, omissions, and lack of coherent inter-document structure (Radev, 2000; Hendrickx et al., 2009; Hewapathirana et al., 2023), which are exacerbated when documents are long (Liu et al., 2024a). While there is substantial work on increasing the long-context processing abilities of LLMs (e.g., An et al. (2024b); Gao et al. (2024a); Xiao et al. (2024)), increased context (Lior et al., 2024; Tang et al., 2024) and model size (Wang et al., 2024d) often does not have substantial impacts on MD capabilities.
@@
Continued pre-training of long-context LMs on sets of related documents has proven effective for MD processing (Caciularu et al., 2021; Yasunaga et al., 2022; Caciularu et al., 2023). However, such methods are not extensible to broader tasks and often require large amounts of pre-training data (e.g., Caciularu et al., 2021; Xiao et al., 2022; Caciularu et al., 2023; Peper et al., 2024).
We propose MDCure, a scalable synthetic instruction-following data generation approach to adapt LLMs for MD settings. Inspired by recent work showing the value of small-scale, high-quality supervised data for efficiently improving LMs’ downstream task performance (Zhou et al., 2023; Sachdeva et al., 2024; Tirumala et al., 2024), MDCure addresses the challenges of data scarcity, complexity, and long-context processing in MD tasks. Instead of relying on costly (Cobbe et al., 2021; Yue et al., 2024) and potentially inadequate human-annotated instructions (Srivastava et al., 2023; Chen et al., 2024c) we leverage powerful LLMs such as GPT-4 (OpenAI et al., 2024) to generate and filter synthetic instructions.
MDCure automatically curates high-quality MD instruction data to enhance LLMs’ MD capabilities. It divides the process into two phases: Generation, where zero-shot prompt templates are used to create complex, cross-text instructions from a set of related documents provided as input context, and Filtering, where a multi-objective, MD-specific reward model, MDCureRM, assesses the generated instructions based on key criteria. This filtering ensures high-quality, diverse, and MD-focused instructions. We find that MDCureRM is pivotal to improving the constitution of the resulting instruction dataset. Instead of generating documents synthetically, by grounding the instructions in real-world texts, MDCure produces realistic and relevant instruction data, improving the model’s ability to synthesize information across multiple sources.
@@
We showcase the efficacy of MDCure through extensive experiments on 6 multi-document benchmarks in over 6 content domains. To analyze the impact of MDCure at scale, we create MDCure instruction datasets of size 12K, 36K, and 72K to fine-tune FlanT5 (Chung et al., 2022), Qwen2-Instruct (Yang et al., 2024), and LLAMA3.1-Instruct (Dubey et al., 2024) models up to 70B parameters in size. Our results show that despite its simplicity, MDCure enables models to consistently achieve superior performance versus existing MD pre-training approaches and outperform prior instruction-following LLMs in both zero- and few-shot settings, with gains of up to 75.5% average improvement. Our contributions are as follows:
  • MDCure, a scalable and general procedure for creating high-quality multi-document instructions that target cross-document relationships and are grounded in real-world contexts, which is robust across model sizes and types.
  • A suite of instruction data complementary to collections such as FLAN (Longpre et al., 2023) for improving MD task performance.
  • A novel set of criteria by which to effectively assess the utility of instruction samples for improving MD performance.
@@
  • MDCureRM, a reward model used as an evaluator specifically designed for the MD setting, to filter and select high-quality MD instruction data in a cost-effective way.
@@ 2 Related Work
Multi-Document Modeling. A common approach for MD modeling with LLMs is to perform flat concatenation of all input documents to process with a long-context model (e.g., Longformer (Beltagy et al., 2020), LongT5 (Guo et al., 2022), CoLT5 (Ainslie et al., 2023)). However, such models lack sufficient skill at synthesizing cross-document information (Caciularu et al., 2022; Wolhandler et al., 2022; Chen et al., 2023; Hosseini et al., 2024). More recently, pre-training-based approaches have proven valuable and involve (continual) training of long-context LMs on instances formulated over sets of related documents. Methods such as token unmasking and masked sentence generation have been used to enhance cross-document information utilization for summarization and classification tasks (Zhang et al., 2020; Caciularu et al., 2021; Yasunaga et al., 2022; Xiao et al., 2022). QAMDen (Caciularu et al., 2023) extends these works with a pre-training objective focused on cross-document question-answering. PELMS (Peper et al., 2024) subsequently improves MDS performance by using semantic clustering to determine salience and intentional sentence removal to boost abstractiveness. Other notable recent works leverage LLMs to better perform targeted content selection for MDS and QMDS (Kurisinkel and chen, 2023; Kurisinkel and Chen, 2023; Li et al., 2023a). In contrast to such methods, our approach employs synthetic instruction generation to overcome limitations of heuristic-based data generation and the computational expense of pre-training, offering a scalable way to adapt LLMs for a wide variety of MD tasks.
@@
Synthetic Instruction Generation. Instruction tuning (IT) is a commonly used approach to align pre-trained LMs for various tasks by training on supervised instruction-answer pairs (Sanh et al., 2022; Wei et al., 2022). However, the scarcity of domain-specific instruction data remains a challenge, as manual annotations are costly and impractical at scale (Song et al., 2024; Ziegler et al., 2024). To address this, recent works have focused on leveraging LLMs for synthetic data generation (Honovich et al., 2023; Wang et al., 2023; Chen et al., 2024a). Most of these efforts target general single-document tasks, but some, such as Chen et al. (2024c), delve into MHQA by generating and merging single-hop QA pairs. While effective, these approaches often focus on MHQA and rely on complex frameworks. Similarly, Lupidi et al. (2024) generates synthetic MHQA data but is limited to two-hop tasks. In contrast, MDCure targets broader MD tasks, with varying complexity, providing a simpler, more scalable solution. Data filtering is also critical, as synthetic data can be noisy or nonfactual (Yue et al., 2024; Ziegler et al., 2024). Several works employ LLMs for selecting quality data at scale (Du et al., 2023; Li et al., 2024), with fine-grained rewards used to enhance task-specific filtering (Wang et al., 2024c). Inspired by these approaches, we design MDCureRM, a novel multi-objective MD-specific reward model to score the diversity and quality of MD instruction data, which we use to refine initially generated instruction pools.
@@ 3 MDCure
MDCure (Fig. 1) utilizes a two-step approach to produce high-quality MD instruction samples over sets of related documents. These samples can be used via instruction tuning to improve base LLMs’ MD modeling capabilities.
@@ 3.1 Generation Phase
In the Generation phase, we sample from a set of carefully tailored zero-shot templates (App. A.2) to prompt a black-box LLM to generate candidate instructions for a given set of documents. The aim is to produce relevant instruction-answer pairs concerning information spanning multiple documents in the context of the LLM. Versus Caciularu et al. (2023), our instructions are more diverse and open-ended. These prompt templates are designed to encourage cross-document reasoning by requiring answers that synthesize information from multiple documents or text segments. They encourage the use of multiple sources instead of only one document, thus involving deeper connections among the provided texts. The wide variety of templates ensures diverse task formulations, ranging from yes/no answers to detailed summaries, which together help generate synthetic MD data that reflects real-world complexity and encourages LLMs to develop strong cross-document comprehension. We analyze the effect of different prompt approaches on the quality of resulting generations in §5.4.
@@
Dataset Source Selection. MDCure’s generation focuses on constructing candidate instructions from sets of documents, which may be sourced from existing datasets of related texts or created by clustering general text corpora. In our experiments, we use topically related document sets from the news domain to model cross-text relationships, building on prior work that shows training with related document sequences can enhance model performance (Caciularu et al., 2021, 2023; Shi et al., 2023). Nonetheless, our generation procedure is equally functional over sets of unrelated documents, which may be utilized to target creation of instruction samples concerning processing of conflicting or irrelevant cross-document information. We base our generated instructions on the NewSHead dataset (Gu et al., 2020). Additional details are in App. A.1.
@@
Generator Model. MDCure is not generally dependent on any specific generator LLM. In combination with our proposed filtering procedure (§3.2), any strong instruction-following LLM can be used to construct effective instruction data. We utilize GPT-3.5-Turbo in our experiments to balance generation quality and costs. [^1] Other cost efficient LLMs such as GPT-4o-mini or Gemini 1.5 Flash can also be used; our generation pipeline was done before the release of these more cost-efficient models. Additionally, preliminary experiments with open-source generators such as LLAMA2-13B led to weak candidate pools that worsened instruction-tuned models’ MD performance.
Instruction Format. We format the MD instruction inputs to place the generated instruction after the source documents, but alternate formats may be equally effective. Since the expected output length of downstream MD tasks has a high variance (e.g., keyword based QA vs. long-form summarization) we include a brief direction indicating the expected length of the answer (i.e., expected number of words or sentences).222 Our preliminary experiments demonstrated non-negligible performance gains on downstream MD tasks with the attachment of such length direction. This direction is appended to the end of the finalized instruction input, separated by a space (Fig. 3 in App. A.2).
@@
3.2 Filtering Phase
The generation prompts, while carefully designed, can still yield noisy or unsuitable instructions for MD settings. To address this, we apply a model-based filtering phase to remove low-quality, non-factual, or single-document instructions. Inspired by recent work on fine-grained RMs (Wu et al., 2023; Wang et al., 2024b, c), we train a fine-grained, MD-specific reward model, MDCureRM, to evaluate each instruction-answer pair based on six different criteria. These criteria capture both the overall quality of the instruction-response pairs and their effectiveness in handling multi-document content and are depicted in Fig. 2. This fine-grained RM-as-a-Judge mechanism refines the final curated instruction set to yield high-quality training data.
Data. Training data for MDCureRM is obtained using an LLM-based pipeline without requiring human annotation. We first use GPT-4o-mini (OpenAI et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) to generate MD instruction-answer pairs of varying quality. [^3] We select these models due to their strong instruction-following capabilities, while considering the generation costs. This results in roughly 20,000 data points. We then prompt GPT-4o to assign scores for each sample according to the six criteria, to be utilized as annotations to train the reward model. The specific prompt used to do so can be found in App. A.5. The target reward scores are normalized to the range $[0,1]$. Although using GPT-4o directly as the RM is also possible, it is prohibitively costly to scale. Training MDCureRM ensures both cost-effectiveness and scalability for filtering large quantities of long-context data.
Implementation. MDCureRM utilizes the Llama3-8B architecture, with weights initialized from a Bradley-Terry reward model (Xiong et al., 2024) trained from Llama3-8B-Instruct (Dubey et al., 2024). The original output layer is replaced with a linear regression layer, outputting a 6-dimensional rating. The regression head processes the final-layer hidden states. We use Mean Squared Error (MSE) between target and predicted rewards as the loss function, freezing the base model during training, following Wang et al. (2024c).
Data Filtering. During inference, MDCureRM generates a 6-element rating for each candidate instruction. These scores are re-scaled to a 1-5 range and combined in a weighted average (details in App. A.6). The top $N$ highest-scoring samples are selected as the final instructions.

### v3

@@ Abstract
Multi-document (MD) processing is crucial for LLMs to handle real-world tasks such as summarization and question-answering across large sets of documents. While LLMs have improved at processing long inputs, MD contexts still present unique difficulties, including management of inter-document dependencies, redundancy, and incoherent structures. To address this challenge, we introduce MDCure, a scalable and effective instruction data generation framework to enhance the MD capabilities of LLMs without the computational cost of pre-training or reliance on human-annotated data. MDCure generates high-quality synthetic MD instruction data over sets of articles via targeted prompts. We also introduce MDCureRM, a cost-effective, MD-specific reward model to score and filter generated data based on their training utility for MD settings. MDCure is compatible with open- and closed-source models in addition to policy optimization methods such as PPO, enabling even small open-source models to surpass proprietary LLMs as strong generators of high-quality MD instruction data without further data filtering. With MDCure, we fine-tune a wide variety of LLMs up to 70B parameters in size from the FlanT5, Qwen2, and LLAMA3.1 model families. Extensive evaluations on a wide range of MD and long-context benchmarks spanning various tasks and domains show MDCure consistently improves performance over pre-trained baselines and base models by up to 75.1%.
@@
Gabrielle Kaili-May Liu1 Bowen Shi1 Avi Caciularu2 Idan Szpektor2  Arman Cohan1 1Yale University  2Google Research {kaili.liu, arman.cohan}@yale.edu
@@ 1 Introduction
With the rapid expansion of bodies of online texts in domains of science, finance, education, news, and more, multi-document (MD) processing becomes a critical aspect of LLMs’ ability to understand and reason over large quantities of text information, with applications such as (query-focused) multi-document summarization (MDS) (Xiao et al., 2022; Giorgi et al., 2023), multi-hop question-answering (MHQA) (Yang et al., 2018), and cross-document reasoning (Cattan et al., 2021b).
While LLMs can now process tens to hundreds of thousands of tokens, they struggle with the unique demands of multi-document understanding and reasoning: finding and aggregating information across diverse documents, resolving contradictions, handling redundant information, bridging information gaps, and constructing coherent narratives (Hosseini et al., 2024; Lior et al., 2024; Tang et al., 2024; Wang et al., 2024d).
@@
Current approaches to improving MD capabilities primarily rely on continued pre-training of long-context models on related document sets (Caciularu et al., 2021; Yasunaga et al., 2022; Caciularu et al., 2023; Peper et al., 2024). However, such methods are not extensible to broader tasks and require large amounts of pre-training data.
We propose MDCure, an effective and scalable framework for creating high-quality synthetic datasets designed to prioritize MD capabilities of LLMs and adapt LLMs for MD settings. MDCure builds upon recent work showing the value of small-scale, high-quality supervised data for efficiently improving LMs’ downstream task performance (Zhou et al., 2023; Sachdeva et al., 2024; Tirumala et al., 2024), in addition to addressing the challenges of data scarcity and complexity in MD tasks. While curating organic, human-created training datasets might seem plausible, such approaches are often prohibitively expensive (Cobbe et al., 2021a; Yue et al., 2024) and suffer from limited scope, failing to capture complex reasoning patterns involved in MD settings (Srivastava et al., 2023; Chen et al., 2024c). Synthetic data leveraging frontier LLMs is becoming increasingly essential in post-training of open-source models Abdin et al. (2024); Lambert et al. (2024). Our proposed MDCure is the first framework for creation of high-quality and MD-focused post-training datasets. Not only does MDCure achieve strong cross-task and cross-domain generalization in long-context MD settings, it also offers practical value due to its compatibility with both cost-effective commercial LLMs (e.g., GPT-3.5) and moderately-sized open-source LLMs (e.g., LLAMA3.1-70B).
MDCure divides the process to automatically curate high-quality MD instruction data into two phases. During Generation, zero-shot prompt templates are used to create complex, cross-text instruction-following prompt-response pairs from a set of related documents provided as input context. During Filtering, a multi-objective, MD-specific reward model, MDCureRM, is trained to assess the generated instructions based on targeted criteria to ensure high-quality, diverse, MD-focused samples. We find that MDCureRM is pivotal to improving the constitution of the resulting instruction dataset. Versus frontier LLMs, MDCureRM overcomes limitations on the tradeoff between cost and quality, offering greater enhancements to MD data quality while also reducing filtering costs. Moreover, as we show, MDCureRM seamlessly integrates with policy optimization methods such as PPO to iteratively improve synthetic MD instruction data generation policies, enabling even small open-source models such as LLAMA3.1-8B-Instruct to surpass SOTA proprietary models as strong generators of high-quality MD instruction data without further filtering or data curation.
@@
We showcase the effectiveness of MDCure through extensive experiments on 6 MD and long-context benchmarks in over 6 content domains. To analyze the impact of MDCure at scale, we create MDCure instruction datasets of size 12K, 36K, and 72K to fine-tune FlanT5 (Chung et al., 2022), Qwen2-Instruct (Yang et al., 2024), and LLAMA3.1-Instruct (Dubey et al., 2024) models up to 70B parameters in size. Our results show that MDCure enables models to consistently achieve superior performance versus existing MD pre-training approaches, and to generalize well across diverse tasks and texts. Further, LLMs post-trained using our MDCure datasets consistently outperform prior instruction-following LLMs and strong long-context baselines, in both zero- and few-shot settings, with gains of up to 75.1% average improvement. Our contributions are as follows:
  • MDCure, a novel framework for obtaining high-quality, MD-focused post-training datasets to improve MD instruction-following capabilities of LLMs, which achieves strong results with both open-source and proprietary cost-effective LLMs.
@@
  • MDCureRM, a novel evaluator reward model specifically designed for the MD setting, to quality-filter MD instruction data more successfully and inexpensively versus proprietary LLMs.
  • A suite of instruction data complementary to collections such as FLAN (Longpre et al., 2023) for improving MD task performance of any LLM. [^1] We release our code at https://github.com/yale-nlp/MDCure.
@@ 2 Related Work
Multi-Document Modeling. A common approach for MD modeling with LLMs is to perform flat concatenation of all input documents to process with a long-context model (e.g., Longformer (Beltagy et al., 2020), CoLT5 (Ainslie et al., 2023)). However, such models lack sufficient skill at synthesizing cross-document information (Caciularu et al., 2022; Wolhandler et al., 2022; Chen et al., 2023; Hosseini et al., 2024). More recently, pre-training-based approaches have proven valuable and involve (continual) pre-training of long-context LMs on instances formulated over sets of related documents using masking and infilling objectives (Zhang et al., 2020; Caciularu et al., 2021; Xiao et al., 2022; Yasunaga et al., 2022). Other notable recent works leverage LLMs to better perform targeted content selection for MDS and QMDS (Caciularu et al., 2023; Kurisinkel and chen, 2023; Kurisinkel and Chen, 2023; Li et al., 2023a; Peper et al., 2024). In contrast, our MDCure is, to our knowledge, the first systematic framework for high-quality MD-specific instruction generation with LLMs to overcome the computational expense of pre-training, the limitations of heuristic-based data generation, and limited generalization of MD performance.
@@
Synthetic Instruction Generation. Instruction tuning (IT) aligns LMs for diverse tasks via supervised instruction-answer pairs but is hindered by (1) the difficulty of obtaining high-quality instructions (Song et al., 2024; Ziegler et al., 2024), and (2) unavailability of high-quality open-source instruction data Lambert et al. (2024). To address this, recent works have focused on leveraging proprietary LLMs for synthetic data generation (Honovich et al., 2023; Wang et al., 2023; Chen et al., 2024a). Most of these efforts target general single-document tasks. A few, such as Chen et al. (2024c); Lupidi et al. (2024), target MHQA by generating and merging single-hop and two-hop QA pairs. While effective, these approaches focus on QA without generalizing to broader MD tasks and rely on complex frameworks. In contrast, MDCure targets MD tasks broadly and enables cross-task generalization, while providing a simpler, more scalable solution to generate complex MD instruction data. As synthetic data can be noisy, data selection using LLMs can be useful in filtering low quality data (Yue et al., 2024; Ziegler et al., 2024). Recent approaches also leverage fine-grained rewards for task-specific data selection (Wang et al., 2024c). Building on this, we introduce MDCureRM, a multi-objective MD-specific reward model that refines instruction data effectively and cost-efficiently and which serves well both as a standalone judge and in RLAIF (Bai et al., 2022; Lee et al., 2024) settings. Overall, with our MDCure framework, we construct and release open-source, high-quality instruction-following datasets that enables significant improvements on MD tasks. As we show, MDCure is not dependent on any specific LLM, works well with both open and proprietary models, and generalizes to many tasks.
@@ 3 MDCure
MDCure (Fig. 1) utilizes a two-step approach to produce high-quality MD instruction samples over sets of related documents. These samples can be used via IT to improve base LLMs’ MD abilities.
@@ 3.1 Generation Phase
During Generation, MDCure performs targeted construction of candidate instructions for a given set of documents by sampling from a set of carefully tailored zero-shot templates (App. B.2) to prompt a generator LLM. Adopting a principled approach, we design our prompt templates to encourage cross-document reasoning by requiring answers that synthesize information from multiple documents, with wide template variety ensuring diverse task formulations, ranging from single-word answers to detailed summaries. This yields synthetic MD data that reflects real-world complexity and encourages LLMs to leverage deep connections among texts to strengthen cross-document comprehension skills. Versus Caciularu et al. (2023), our instructions are more diverse and open-ended. Qualitative examples and diversity analysis of our MDCure instructions [^2] Details regarding instruction format are in App. B.4. are provided in App.s B.5 and B.6, respectively. A systematic analysis of the effect of different prompt approaches on the quality of resulting generations is provided in §A.1.
@@
Dataset Source Selection. MDCure’s generation focuses on constructing candidate instructions from sets of documents, which may be sourced from existing datasets of related texts or created by clustering general text corpora. In our experiments, we use topically related document sets from the news domain to model cross-text relationships, building on prior work that shows training with related document sequences can enhance model performance (Caciularu et al., 2021, 2023; Shi et al., 2023). Nonetheless, our generation procedure is equally functional over sets of unrelated documents, which may be utilized to target creation of instruction samples concerning processing of conflicting or irrelevant cross-document information. We base our generated instructions on the NewSHead dataset (Gu et al., 2020). Additional details are in App. B.1.
@@
Generator Model. MDCure is not generally dependent on any specific generator LLM. In combination with our proposed filtering procedure (§3.2), any strong instruction-following LLM can be used to construct effective instruction data. We utilize GPT-3.5-Turbo in experiments to balance generation quality and cost. We also show the compatibility and comparable efficacy of MDCure with use of open-source generator LLAMA3.1-70B in §5.4.333Other cost efficient LLMs such as GPT-4o-mini or Gemini 1.5 Flash and open-source LLMs such as LLAMA3.3-70B can also be used; our generation pipeline was done before the release of these more cost-efficient models.
3.2 Filtering Phase
The generation prompts, while carefully designed, can still yield noisy or unsuitable instructions for MD settings. To address this, we apply a model-based filtering phase to remove low-quality, non-factual, or single-document instructions. Inspired by recent work on fine-grained RMs (Wu et al., 2023; Wang et al., 2024b, c), we train a fine-grained, MD-specific reward model, MDCureRM, to evaluate each instruction-answer pair based on six different criteria. These criteria capture both the overall quality of the instruction-response pairs and their effectiveness in handling multi-document content and are depicted in Fig. 2. This fine-grained RM-as-a-Judge mechanism refines the final curated instruction set to yield high-quality training data. We analyze the criticality of each MDCureRM scoring criterion and the effect of their relative weightings via ablation studies in App. A.2. A meta-evaluation to analyze the alignment of MDCureRM with human annotations versus GPT-3.5-Turbo is provided in App. E, where we observe agreements in line with LLM evaluation literature (Liu et al., 2024a). §5.5 further explores the efficacy of using reward signals provided by MDCureRM to train a custom MD instruction generator model via PPO without the need for post-generation filtering.
@@ 3.2 Filtering Phase
Data. Training data for MDCureRM is obtained using an LLM-based pipeline without requiring human annotation. We first use GPT-4o-mini (OpenAI et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) to generate MD instruction-answer pairs of varying quality. [^4] We select these models due to their strong instruction-following capabilities, while considering the generation costs. This results in roughly 20,000 data points. We then prompt GPT-4o to assign scores for each sample according to the six criteria, to be utilized as annotations to train the reward model. The specific prompt used to do so can be found in App. B.8. The target reward scores are normalized to the range $[0,1]$. Although using GPT-4o directly as the RM is also possible, it is prohibitively costly to scale. Training MDCureRM ensures both cost-effectiveness and scalability for filtering large quantities of long-context data.
Implementation. MDCureRM utilizes the Llama3-8B architecture, with weights initialized from a Bradley-Terry reward model (Xiong et al., 2024) trained from Llama3-8B-Instruct (Dubey et al., 2024). The original output layer is replaced with a linear regression layer, outputting a 6-dimensional rating. The regression head processes the final-layer hidden states. We use Mean Squared Error (MSE) between target and predicted rewards as the loss function, freezing the base model during training, following Wang et al. (2024c). We discuss alternative architectures and the design of MDCureRM in App. C.2.
Data Filtering. During inference, MDCureRM generates a 6-element rating for each candidate instruction. These scores are re-scaled to a 1-5 range and combined in a weighted average (details in App. B.9). The top $N$ highest-scoring samples are selected as the final instructions.